how would I graph y2=12x (is this a parabola?)
Answer by stanbon(75887)

how would I graph y^2=12x (is this a parabola?)
It is a parabola opening to the right.
---------
Solve for "x":
x = (1/12)y^2
If y = 0, x=0
If y = +sqrt(12), x = 1
If y = -sqrt(12), x = 1
So the vertex is at (0,0)
The other two points are symmetric to the x-axis and to the right.
Continue the graph as a parabola opening to the right.
==============
cheers,
stan H.